OpenStudy (anonymous):

Double check my work? Is the answer to (x + 6)(x − 6) and (x − 6)(x − 6) = x^2 - 36 and x^2 + 36?

OpenStudy (anonymous):

Yes you're correct :)

OpenStudy (anonymous):

The first one is right, the second should be x^2 -12x +36

OpenStudy (anonymous):

no wait, on the 2nd one it's x^2-12x+36, first one is right.

OpenStudy (anonymous):

Oh, what makes it that way? If you don't mind explaining (:

OpenStudy (cwrw238):

(x - 6)(x - 6) use the distributive law first multiply the contents of second parentheses by x, then by -6: = x*x + x*(-6) - 6(x) -6(-6) = x^2 -6x - 6x + 36 the 2 middle terms simplify: = x^2 - 12x + 36

OpenStudy (cwrw238):

if you do same thing with (x + 6)(x - 6) the 2 middle terms are -6x + 6x which = 0 so they disappear last term -s -6 * 6 = -36 giving x^2 - 36

OpenStudy (anonymous):

thank you all!

OpenStudy (cwrw238):

yw